a parrticle A is projected vewrtically upwards . another particle B of same mass is prijected at an angle 45 .both reach at same height . the ratio of intial kinetic energy of A to B ?

Grade:

2 Answers

Puneet Mittal AskiitiansExpert-IITD
13 years ago

Hi Srishty,

Let Initial speeds of A be Va and of B be Vb.

let height be h.

Thus 2gh = (Va)    and  (Vb(sin45))2 = 2gh

Thus, (Va)2 = (Vb)2*(1/2)

Thus, (Va)2/(Vb)2  = 1/2

The ratio of Initial Kinetic energies of A and B is 1/2 because they have same Mass.

Please feel free to ask as many questions you have.

Puneet

Karthik Eyan
13 years ago

For a 1 D motion,

H max = u^2/2g

In an oblique projection, H max = v^2 sin^2 (angle) / 2g

Here heights are same.

So :u^2 = v^2 sin^2(45)

u^2 / v^2 = 1

So, ratio of internal KE is 1:1 as masses are same........